OpenStudy (anonymous):

Find the domain of the composite function F(f(x)). F(x)= x+3, g(y)=4/ sqr root of y G(f(x))= 4/ sqr root of x+3

OpenStudy (anonymous):

when the function is under a square root, the function's domain is everywhere that makes the function non-negative, also since there is a fraction you cannot have a zero in the denominator

OpenStudy (anonymous):

Denominator cannot be 0. For what value of x does sqrt(x+3) = 0?

OpenStudy (anonymous):

oops right. The expression in the radical sign x+3 must be greater than zero. AND it cannot equal zero. So domain is x+3\[\ge\]0. Solve that inequality for x.

OpenStudy (anonymous):

Oops again I mean >

OpenStudy (anonymous):

so looking at this closely, sqrt(x+3) you are look for an x that doesnt result in a negative so x must be greater than or equal to -3, but it is also under a fraction so when it is equal to -3 it becomes zero and we do not want that in the denominator so its everywhere greater than -3 but not equal to 3 or in other words\[(-3,\infty)\]

OpenStudy (anonymous):

i mean -3 on that last line not 3